# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ H.1 The Kaplan-Meier estimator
$S_{E}(t)=\text{Pr}(E\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(E=t_{i})\right)\text{Pr}(E\geq t_{i})$ (H.1)
Second, suppose that $E$ and $C$ are independent. Then, we can write
$S_{E}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(E=t_{i},C\geq t_{i})\right)\text{Pr}(E\geq t_{i},C\geq t_{i})$ (H.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(E,C)=t_{i},E\leq C)\right)\text{Pr}(\min(E,C)\geq t_{i})$ (H.3)
@@
$S_{E}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (H.4)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{1}{n_{i}}\right)$ (H.5)
H.2 The Copula-Graphic estimator
This derivation is reproduced from [Emura and Chen, 2018, Ch. 4].
@@ H.2 The Copula-Graphic estimator
$P(E>t,C>c)=\phi_{\theta}^{-1}\big{[}\phi_{\theta}\{S_{E}(t)\}+\phi_{\theta}\{S_{C}(t)\}\big{]},$ (H.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3], and $S_{E}(t)=\text{Pr}(E>t)$ and $S_{C}(t)=\text{Pr}(C>t)$ are the marginal survival functions. Assume that $S_{E}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{E}(t_{i})\neq S_{E}(t_{i}-dt)$ and $S_{C}(t_{i})=S_{C}(t_{i}-dt)$. Setting $t=t_{i}$ in Equation H.6, we have
$\text{Pr}(E\geq t_{i},C\geq t_{i})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}$ (H.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(E>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}\mathbbm{1}[t_{\ell}>t_{i}]$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}(\frac{n_{i}-1}{n})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (H.8)
We set $t=c=t_{i}-dt$ in Equation H.6 and then estimate $\text{Pr}(E>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{E}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\delta_{i}=1$ (H.9)
Equations H.8 and H.9 result in the system of difference equations
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\delta_{i}=1$ (H.10)
We apply the standard constraint that $S_{E}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{E}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{E}(t_{i})\}-\phi_{\delta}\{S_{E}(t_{i}-dt)\}]$ (H.11)
$=\sum_{t_{i}\leq t,\delta_{i}=1}\phi_{\delta}\big{(}\frac{n_{i}-1}{n}\big{)}-\phi_{\delta}(\frac{n_{i}}{n})$ (H.12)
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-\log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [Clayton, 1978], the CG estimator is written as
$S_{E}(t)=\Bigl{[}1+\sum_{t_{i}\leq t,\delta_{i}=1}\Bigl{\{}(\frac{n_{i}-1}{n})^{-\theta}-(\frac{n_{i}}{n})^{-\theta}\Bigr{\}}\Bigr{]}^{-1/\theta}$ (H.13)

### v3

@@ G.1 The Kaplan-Meier estimator
$S_{T}(t)=\text{Pr}(T\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\text{Pr}(T=t_{i})}{\text{Pr}(T\geq t_{i})}\right).$ (G.1)
Second, suppose that $T$ and $C$ are independent. Then, we can write
$S_{T}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i},C\geq t_{i})\right)\text{Pr}(T\geq t_{i},C\geq t_{i})$ (G.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(T,C)=t_{i},T\leq C)\right)\text{Pr}(\min(T,C)\geq t_{i}).$ (G.3)
@@
$S_{T}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (G.4)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{1}{n_{i}}\right).$ (G.5)
G.2 The Copula-Graphic estimator
This derivation is reproduced from [4, Ch. 4].
@@ G.2 The Copula-Graphic estimator
$P(T>t,C>t)=\phi_{\theta}^{-1}\big{[}\phi_{\theta}\{S_{T}(t)\}+\phi_{\theta}\{S_{C}(t)\}\big{]},$ (G.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3], and $S_{T}(t)=\text{Pr}(T>t)$ and $S_{C}(t)=\text{Pr}(C>t)$ are the marginal survival functions. Assume that $S_{T}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{T}(t_{i})\neq S_{T}(t_{i}-dt)$ and $S_{C}(t_{i})=S_{C}(t_{i}-dt)$. Setting $t=t_{i}$ in Equation G.6, we have
$\text{Pr}(T\geq t_{i},C\geq t_{i})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (G.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(T>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}\mathbbm{1}[t_{\ell}>t_{i}]$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}\left(\frac{n_{i}-1}{n}\right)=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (G.8)
We set $t=c=t_{i}-dt$ in Equation G.6 and then estimate $\text{Pr}(T>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\quad\delta_{i}=1$ (G.9)
Equations G.8 and G.9 result in the system of difference equations
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\quad\delta_{i}=1$ (G.10)
We apply the standard constraint that $S_{T}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{T}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{T}(t_{i})\}-\phi_{\delta}\{S_{T}(t_{i}-dt)\}]$ (G.11)
$=\sum_{t_{i}\leq t,\delta_{i}=1}\phi_{\delta}\left(\frac{n_{i}-1}{n}\right)-\phi_{\delta}\left(\frac{n_{i}}{n}\right),$ (G.12)
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-\log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [10]
$S_{T}(t)=\left[1+\sum_{t_{i}\leq t,\delta_{i}=1}\left\{\left(\frac{n_{i}-1}{n}\right)^{-\theta}-\left(\frac{n_{i}}{n}\right)^{-\theta}\right\}\right]^{-1/\theta}.$ (G.13)